From mar@roe.ac.uk Sun Nov 12 18:07:12 2017 Date: Sun, 12 Nov 2017 18:07:05 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , dmr@roe.ac.uk, Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Roy Williams , Stelios Voutsinas Subject: WFAU data services meeting minutes, 1 November 2017 WFAU data services meeting minutes, 1 November 2017 Present: RPB, MAR, RSC, RDW, RGM, MSH Apologies: DMR, STV, NCH, KTN, AL, ACD, NJC, ETWS DoNM: 11:00 am Wednesday 15th November 2017, Villas Meeting Room MAR in the chair (room is booked). Actions discharged this time: ----------------------------- None Actions carried forward from 19/10/2017 meeting: ---------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan DR13 is installed but not yet cross-mounted. However, we still need to decide which SDSS releases to keep so the present action should remain. ACTION: MAR to write some H-ATLAS documentation once H-ATLAS is available. Continues. ACTION: NJC to check if any of the new VISTA surveys warranted a P98 release. Continues ACTION: RSC/NJC look into ppErrBit inconsistencies. Continues ACTION: ALL to consider advantages/disadvantages of combining or keeping VVV/VVVX separate. Continues, NJC had had some discussions with the PIs, but no conclusion yet. Actions carried forward from the Backup Strategy Meeting held on 27/1/2017: -------------------------------------------------------------------------- ACTION: RPB to create a twiki page documenting the backup procedure. These notes should be usable by someone not routinely involved in the backups and, for example, include the physical location of the tapes. Tuesday 28 February 2017 was subsequently agreed as a target date for completing these notes. A stub page for RPB to complete is available at: http://apache.roe.ac.uk/twiki/bin/view/WFAU/BackupsProcedures Continues. ACTION: RPB to create a twiki page documenting the procedure to restore datasets from a backup. These notes should be usable by someone not routinely involved in the backups and, for example, include the physical location of the tapes. A stub page for RPB to complete is available at: http://apache.roe.ac.uk/twiki/bin/view/WFAU/BackupsRecovery Continues. ACTION: RGM to ask Steven Duffield to review the WFAU backup procedures. Continues. This action could not progress until RPB had written notes describing the backup and recovery procedures. ACTION (amended): MSH (was all) set up a meeting to review tests of SQL server on Linux prior to the next telecon with John Hopkins. Continues, awaits results of testing. Actions carried forward from the SQL Server on Linux Meeting 27/1/2017: --------------------------------------------------------------------------- ACTION: NJC to run the "testall" script as a basic sanity check. Check all file paths and change to Linux style. Should run in 30 minutes. This failed fairly quickly as XP_CMDSHELL not available under Linux. NEW ACTION: RPB to check out how file permissions operate under MSSQL on Linux with a view to replacing XP_CMDSHELL with ssh operations. ACTION: RPB to run curation scripts on the WSA database to test re-doing a previous release. ACTION: MSH to check the crossmount/crossmatch/crossshare capabilities, connect two more servers, one Windows and one Linux. Candidate machines are ramses1 and ramses2. Continues. Specific points and new actions: -------------------------------- Project management: * RGM reported that Edinburgh's e-Infrastructure bid had been successful. This would provide 1Pb of storage and cloud compute. It was expected that a VVV release would be used as a testbed. * RGM would shortly be arranging the next WFAU Quarterly. WFCAM, VISTA and VST updates: * Nothing for the minutes Networking: * All WFAU networks were working nominally. WSA/VSA/OSA/GES Operations: * VSA P98 releases were proceeding. * Constraints were still being applied to VVV. * There had been issues with the astroAnalysis content of the GESiDR5 release that ETWS had been trying to address whilst on leave. * RPB noted that CU19 of the VMC P98 release had initially fallen over claiming the SOURCE_FG was full. There was plenty of room on disk and the FG should automatically grow, so it was not clear why the error was thrown. A manual FG size increase fixed the issue. ACTION: RPB to add note to twiki to note filegroup issue and fix. * Investigations were still ongoing on the UKIDSS field groupings and deep stacks. Hardware and Systems: * A purchase order had gone in for the new DB node. * MSH noted that there had been an increase in disk failures (eg wahkahre) since the power outage. * MSH had been working on automatic shutdown scripts for our machines in the event of another power failure. This was working well on the Linux curation servers and windows servers but some of the NAS boxes were not able to be managed this way. However as other ROE/ATC machines would continue to operate on the UPS whilst the AC was off the temp in C2 would still rise and eventually cause problems for disks on our shutdown machines. ACTION: MSH to email RGM and Jim a short summary of the current situation with regard to C2 and a significant power outage. * MSH reported that a Linux upgrade of our curation servers was required. He would wait for ETWS's return before proceeding. Status of Backups: * RPB reported that backups were running well. Review of Software Release Deadlines: * The prioritised list of software items is available at: http://apache.roe.ac.uk/twiki/bin/view/WFAU/PrepareSoft Other Software: Nothing for the minutes. Review of survey deadlines: * Progress towards forthcoming survey deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * VMCv20171101 had been released. Non-survey Data Release: * MAR asked RPB to try and get a few releases running. Astrogrid deployment, VO & Data Analysis services: * There had been issues registering the OSA TAP service. Miscellaneous: Nothing for the minutes